The resistance π‘…πœƒ of a particular resistance thermometer at a Celsius temperature πœƒ as measured by a constant-volume gas thermometer is given by

π‘…πœƒ = 50.00 + 0.1700πœƒ + 3.00 Γ— 10βˆ’4πœƒ2.

Calculate the temperature as measured on the scale of the resistance thermometer which corresponds to a temperature of 60°𝐢 on the gas thermometer.


1
Expert's answer
2022-04-17T17:06:10-0400

The resistance reading will be the following:


"R(60) = 50.00 +0.1700\cdot 60 + 3.00\times 10^{-4}\cdot 60^2 =61.28"

Answer. "61.28".